2. Comparative advantage. Adam and Boris run a cafeteria. Adam can prepare 1 hamburger in 60 seconds if he works in the kitchen and he can serve 1 customer in 50 seconds whern he works at the checkout. Boris can prepare 1 hamburger in 50 seconds and he can serve 1 customer in 40 seconds. Who has the comparative advantage in which activity? How should they specialize? How many hamburgers they can make and how many customers they can serve a) if they both spend 4 hours preparing hamburgers and 4 hours serving customers, b) if they work 8 hours each and specialize according to their comparative advantages? 3. Substitutes and Complements. Give an example of two complement goods. Wh decreases? at happens if the price of one of them 4. Consumer surplus. The demand for cars is P-800 000-Q The price is 500 000. Compute the consumer surplus. 5. Capital and Labour market. Your nominal salary has increased from CZK 12 000 to CZK 14 000. In the meantime prices went up by 5 %. What is the percent change in the nominal salary? What is the percent change in the real wage? 6. Demand. Estimate the demand function for hotdogs if you know that at the price P-30 people buy 100 hotdogs and at the price P-28 people buy 120 hotdogs. 7. Profit maximisation. You run a McDonalds restaurant Your marginal cost per 1 hamburger is constant at MC-12 The demand for hamburgers is P-30-0,1Q. What is the price that maximizes your profit? 8. Market models. Explain the perfect competition. 9. Explain the Indifference Curve 10. Equilibrium. Find the market equilibrium price and quantity if the demand is P-50-2Q and the supply is P-10+3Q
Answer #1

3. Complementary goods are those goods which jointly satisfy a particular want of consumer. Example; Car and Petrol.

Increase in the price of one good decreases the demand of other. Decrease in the price of one good increases the demand of other. Decrease in the price of petrol increases demand of car and vice-versa.

4.

Price 800,000 C.S 500,000 0 300,000 800,000 Quantity

At price of 500,000; Q = 800,000 - 500,000 = 300,000

Consumer Surplus = 1/2 x base x height = 1/2 x 300,000 x (800,000 - 500,000) = 150,000 x 300,000 = 45000000000

5. % change in nominal salary = (14000 - 12000)/12000 x 100 = 2000/12000 x 100 = 16.67%

% change in price = 5% i.e. inflation is 5%

% change in real wage = % change in nominal wage - % change in price = 16.67% - 5% = 11.67%